Good Afternoon, I am struggling with logarithm problems, and I am required by my instructor to show all the steps on how I obtained my answer. Because I have no idea how to do the problem, I struggle with the steps to follow. If someone could be so kind as to show me exactly step by step what to do, I would be so grateful.
Section 12.3, Pg. 852, # 28
Directions: Use properties of logarithms to condense each logarithmic expression. Write the expression as a single logarithm whose coefficient is 1. Where possible, evaluate logarithmic expressions.
38.) log 250 + log 4
Any help is appreciated.

Answer by nerdybill(7384)

log 250 + log 4
Addition in log can be simplified w/multiplication:
= log(250*4)
= log(1000)
= log(10^3)
= 3